OpenStudy (anonymous):

P(2n+4,3) = 2/3 P(n+4,4)

OpenStudy (triciaal):

what is 4,3?

OpenStudy (anonymous):

btw apperently the answer is 13? oh and we are dealing with P(n,r) stuff or statististics so n=2n+4 and r=3

ganeshie8 (ganeshie8):

(2n+4)(2n+3)(2n+2) = 2/3(n+4)(n+3)(n+2)(n+1)

ganeshie8 (ganeshie8):

cancel cmmon factors and solve n ?

OpenStudy (triciaal):

i had n = 10

OpenStudy (anonymous):

ok give me some time :P gotta solve it to see if i get it

OpenStudy (triciaal):

@ganeshie8 would it work the same as 4C3 and use factorial? then equate and solve for n?

ganeshie8 (ganeshie8):

for 4C3 we need to divide by r! also so thats an extra i think

ganeshie8 (ganeshie8):

but permutation is much simpler : 4P3 = 4x3x2

ganeshie8 (ganeshie8):

nPr = n(n-1)(n-2)...(n-r+1)

ganeshie8 (ganeshie8):

nCr = [ n(n-1)(n-2)...(n-r+1) ] /r!

ganeshie8 (ganeshie8):

so i think combination may give us a messy equation

OpenStudy (triciaal):

thanks

OpenStudy (anonymous):

are we supposed to get 22n^3+86n^2+74n+78=0 at the end?

ganeshie8 (ganeshie8):

(2n+4)(2n+3)(2n+2) = 2/3(n+4)(n+3)(n+2)(n+1) 2(n+2)(2n+3)2(n+1) = 2/3(n+4)(n+3)(n+2)(n+1) 4(2n+3) = 2/3(n+4)(n+3)